A group of hikers is to travel x km by bus at an average speed of 48 km/h to an unknown destination. They then plan to walk back along the same route at an average speed of 4.8 km/h and to arrive back 24 hours after setting out in the bus. If they allow 2 h for lunch and rest, how far must the bus take them?
Answer by mananth(12270)

time taken by bus +time walking + 2hours = 24
...
t= d/r
x/48 + x/4.8 + 2 = 24
x/48 +x/4.8 = 24-2
LCD = 48
(x+10x)/48 = 22
11x=48*22
/11
x= 48*22/11
x=48*2
x=96 km is the distance they have to travel by bus
...
m.ananth@hotmail.ca